OpenStudy (anonymous):

The two spheres above have the same center. One has a radius of 4 cm, and the other has a radius of 5 cm. Approximately how much space is in between the two spheres?

OpenStudy (anonymous):

In order to find the space between, you simply need to subtract the larger volume sphere and the smaller volume sphere. The volume of a sphere is given by: (4/3) * pi * r^3 r is the radius of the sphere For 5 cm: V_5 = (4/3) * pi * 5^3 For 4 cm: V_4 = (4/3) * pi * 4^3 Then subtract: V difference = V_5 - V_4